If 5m3 of a certain oil weighs 4000 kg(f). Calculate the specific weight, mass density and specific gravity of this oil.

1
Expert's answer
2022-02-16T15:05:10-0500

"\gamma=\frac{mg}V=8~\frac{kN}{m^3},"

"\rho=\frac mV=800~\frac{kg}{m^3},"

"SG=\frac{\rho}{\rho_0}=0.8."